a shaft fitted with a fly wheel rotates at 250 rpm and dried a machine the torque of machine varies in a cyclic manner over a period of 3 revolution. The torque rises from 750 Nm to 300Nm uniformly during half revolution and remains constant for the following revolution . If then fall uniformly to 750 Nm during the next half revolution and remains constant for one revolution the cycle being repeated thereafter. Determine the power required to drive the machine and percentage fluctuation in speed if the driving torque applied to the shaft is constant band and mass of the fly wheel is 500 kg with the radius of gyration at 600mm.

1
Expert's answer
2021-12-13T17:08:03-0500

Torque="6π*750+\frac{1}{2}*π(3000-750)+2π(3000-750)+\frac{1}{2}π(3000-750)"

=11250π Nm

"T_{mean}*6\u03c0 Nm=11250\u03c0 Nm"

"T_{mean}=\frac{11250}{6}"

=1875 Nm

P=Tmean*w

=1875*26.2=49125

"\u2206E=mk^2w^2C_s"

"C_s=\frac{8837}{500*0.6^2*26.2^2}"

=0.071